A concrete and rock crusher for demolition work has been purchased for ​$51,000​, and it has an estimated SV of ​$12,000 at the end of its​ five-year life. Engineers have estimated that the following units of production​ (in m3 of crushed​ material) will be contracted over the next five years. Using the units of production depreciation​ method, what is the depreciation allowance in year two​, and what is the BV at the end of year four​?

(please do both parts i will really appreciate it.Thanks!)

EOY 1 2 3 4 5
m^3 14,000 26,000 35,000 15,000 10,000
Answer #1

Given that,

Cost of purchasing concrete & rock crusher = $51,000

Estimated Salvage Value at the end of 5 years = $12,000

EOY 1 2 3 4 5
Production (m^3) 14,000 26,000 35,000 15,000 10,000

Estimated total units of production in 5 years = 14,000 + 26,000 + 35,000 + 15,000 + 10,000 = 100,000 m^3

According to unit of production depreciation method:

Depreciation (per year) = (Original Cost - Salvage Value)/Total Production capacity x Production in the year

Depreciation in Year 1 = (Original Cost - Salvage Value)/Total Production capacity x Production in the year 1 = ($51,000 - $12,000)/100,000 x 14,000 = $5,460

Book Value at the EOY 1 = Original Cost - Depreciation in Year 1 = $51,000 - $5,460 = $45,540

Depreciation in Year 2 = (Original Cost - Salvage Value)/Total Production capacity x Production in the year 2 = ($51,000 - $12,000)/100,000 x 26,000 = $10,140

Depreciation allowance in year 2 = $10,140

Book Value at the EOY 2 = Book Value at the EOY 1 - Depreciation in Year 2 = $45,540 - $10,140 = $35,400

Depreciation in Year 3 = (Original Cost - Salvage Value)/Total Production capacity x Production in the year 3 = ($51,000 - $12,000)/100,000 x 35,000 = $13,650

Book Value at the EOY 3 = Book Value at the EOY 2 - Depreciation in Year 3 = $35,400 - $13,650 = $21,750

Depreciation in Year 4 = (Original Cost - Salvage Value)/Total Production capacity x Production in the year 4 = ($51,000 - $12,000)/100,000 x 15,000 = $5,850

Book Value at the EOY 4 = Book Value at the EOY 3 - Depreciation in Year 4 = $21,750 - $5,850 = $15,900